Determine whether the given number is a solution of the given equation.
1. 35; t+17= 53
2. 36; a-19= 17
Solve using the addition principle.
1. z+9=14

Answer by nabla(475)

1.
35+17=52 which is not equal to 53. (not a solution).
2.
36-19=17 which is equal to 17. (is a solution).


next
1.
z+9=14
z+9-9=14-9
z=5.